Circular No. 6911

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

VARIABLE STAR NEAR NGC 4013 = SUPERNOVA 1989Z? D. Bettoni, Padua Astronomical Observatory, reported her discovery

of a possible supernova (mag ∼ 12) on an unfiltered CCD image obtained of NGC 4013 on May 18.91 UT with the 1.82-m Asiago telescope; the star was confirmed on an image taken on May 20.02. The star lies along the dust lane, ∼ 7′′ east and 8′′ north of the galaxy nucleus, which is very flattened and difficult to define because of the strong dust absorption. The position is very similar to that of the reported SN 1989Z (cf. IAUC 5162), which was reported more than a year after the single observation and was never confirmed. Bettoni adds: “A spectrum taken on May 20 at the Nordic Optical Telescope by J. Hjorth and S. Benetti indicates this to be a star of spectral type A with negligible redshift and no P-Cyg effect. Further examination of images of this galaxy in the literature shows that the star is absent (or very faint) in an image taken in April 1989 with the Lowell 1.1-m telescope (Frey et al., NASA/IPAC Extragalactic Database images), and it is faint (mag ∼ 17) in an image from March 1996 taken with the 2.42-m Hawaii telescope (Tully et al. 1996, A.J. 112, 2471). It is probably a foreground variable star of as-yet-undetermined type. I suspect that this is the same star that was reported by Shaw on IAUC 5162.”

NGC 6814 K. Mukai, Goddard Space Flight Center (GSFC), NASA, and Univer-

sities Space Research Association; G. Madejski, GSFC and University of Maryland; and C. Hellier, University of Keele, write: “We have observed the Seyfert galaxy NGC 6814 with the Rossi X-ray Timing Explorer. On May 18.54 UT, it was detected at a rate of ∼ 15 counts/s in 5 PCUs (subject to background uncertainties of ∼ 5 counts/s), or 3 × 10^{−11} erg cm^{−2} s^{−2} at 2–10 keV. Thus, NGC 6814 appears to be in an unusually high state, an order of magnitude brighter than in 1993 May (1.2 × 10^{−12} erg cm^{−2} s^{−1} at 2–10 keV; George et al. 1998, Ap.J. Supp. 114, 73), and arguably brighter than in any previous x-ray observations. We have also observed the magnetic cataclysmic variable, V1432 Aql (which contaminated previous non-imaging x-ray data on the Seyfert), located 37′ away, and confirmed that cross-contamination in our data is minimal. We will be monitoring both sources at ∼ 4-day intervals until early September

(see http://lheawww.gsfc.nasa.gov/~mukai/v1432.html); coordinated

observations at other wavelengths are strongly encouraged.”

1998 May 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green